\newpage 
\begin{thebibliography}{}

\bibitem{kac_can_10} 
N. Cantarini, V. Kac, {\it Classification of linearly compact simple rigid superalgebras},  Int. Math. Res. Not. (IMRN), {\bf 17} (2010), 3341--3393.

%\bibitem{GR11} 
%M. Goze, E. Remm, {\it $2$-dimensional algebras}, Afr. J. Math. Phys., {\bf 10} (2011), no.~1, 81--91.


\bibitem{perez03}
C. Jiménez-Gestal, J.M. Pérez-Izquierdo,
{\it Ternary derivations of generalized Cayley-Dickson algebras}, 
Comm. Algebra, {\bf 31} (2003), no.~10, 5071--5094.

\bibitem{Kantor72} 
I. Kantor, {\it Certain generalizations of Jordan algebras (Russian)}, 
Trudy Sem. Vektor. Tenzor. Anal., {\bf 16} (1972), 407--499.


\bibitem{Kantor88} 
I. Kantor, 
{\it A universal attracting object in the category of conservative algebras (Russian)}, 
Trudy Sem. Vektor. Tenzor. Anal. No. 23 (1988), 45--48.

\bibitem{Kantor89term} 
I. Kantor, {\it On an extension of a class of Jordan algebras}, 
Algebra and Logic, {\bf 28} (1989), no.~2, 117--121.

\bibitem{Kantor89tril} 
I. Kantor, {\it Terminal trilinear operations}, 
Algebra and Logic, 28 (1989), no.~1, 25--40.

\bibitem{Kantor89trudy}
I. Kantor, {\it Some problems in $L$-functor theory (Russian)}, 
Trudy Inst. Mat. (Novosibirsk), no.~16 (1989), Issled. po Teor. Kolets i Algebr, 54--75.

\bibitem{Kantor90}
I. Kantor, {\it An universal conservative algebra}, 
Siberian Math. J., {\bf 31} (1990), no.~3, 388--395.


\bibitem{kay13}
I. Kaygorodov, {\it On $(n+1)$-ary derivations of simple $n$-ary Malcev algebras}, 
St. Petersburg Math. J., {\bf 25} (2014), no.~4, 575--585.

\bibitem{kay14}
I. Kaygorodov, {\it $(n+1)$-ary Derivations of Semisimple Filippov algebras}, 
Math. Notes, {\bf 96} (2014), no.~2, 208--216.

\bibitem{kay} 
I. Kaygorodov, {\it   On the Kantor Product},  
arXiv:1506.00736

\bibitem{kaylopo}  
I. Kaygorodov, A. Lopatin, Yu. Popov, {\it Conservative algebras of 2-dimensional algebras},  
arXiv:1411.7744


\bibitem{GR11} 
H. Petersson, {\it The classification of two-dimensional nonassociative algebras}, 
Results Math., {\bf 37} (2000), no. 1-2, 120--154. 

\bibitem{shestakov14}
A. Shestakov, 
{\it Ternary derivations of Jordan superalgebras}, 
Algebra and Logic, {\bf 53} (2014), no.~4, 323--348. 

\end{thebibliography}

\documentclass[12pt,oneside,reqno]{article}
\usepackage[utf8]{inputenc}
\usepackage[english]{babel}
\usepackage{amsmath}
\usepackage{amssymb}
\usepackage{amsfonts}
\usepackage{mathrsfs}
\newtheorem{Th}{Theorem}
\newtheorem{Lem}[Th]{Lemma}

\newtheorem{Remark}[Th]{Remark}

\newtheorem{corollary}[Th]{Corollary}

\newcommand{\FF}{{\mathbb{F}}} 

\newcommand{\Aut}{{\mathrm{Aut}}}

\newcommand{\Id}{{\rm Id}}
\newcommand{\OO}{\mathcal{O}}
\newcommand{\SSS}{\mathcal{S}}
\newcommand{\TT}{\mathcal{T}}

\begin{document}

\sloppy


	{\bf Conservative algebras of $2$-dimensional algebras, II}

	\
	
	Ivan Kaygorodov$^{a,b}$, Yury Volkov$^{c,d}$ \\

	\medskip

{\tiny
$^{a}$ Universidade Federal do ABC, CMCC, Santo Andr\'{e}, Brazil.

$^{b}$ Sobolev Institute of Mathematics,  Novosibirsk, Russia.

$^{c}$ Universidade de S\~{a}o Paulo, IME, Sao Paulo, Brazil.

$^{d}$ Saint Petersburg state university, Saint Petersburg, Russia.
\smallskip

    E-mail addresses:

    Ivan Kaygorodov (kaygorodov.ivan@gmail.com),
    
    Yury Volkov (wolf86\_666@list.ru).

}

\

\medskip

{\bf Abstract.} In 1990 Kantor defined the conservative algebra $W(n)$ of all algebras (i.e. bilinear maps) on the $n$-dimensional vector space. If $n>1$, then the algebra $W(n)$  does not belong to any well-known class of algebras (such as associative, Lie, Jordan, or Leibniz algebras). 
We describe automorphisms, one-sided ideals, and idempotents of $W(2).$ 
 Also similar problems are solved for the algebra $W_2$ of all commutative algebras on the 2-dimensional vector space  and for the algebra $S_2$ of all commutative algebras with trace zero multiplication on the 2-dimensional vector space.

\medskip

{\bf Keywords:} bilinear maps, conservative algebra, ideal, automorphism, idempotent.

{\bf 2010 MSC: } 15A03; 15A69; 17A30; 17A36

\medskip

\section{Introduction}
During this paper $\FF$ is some fixed field of zero characteristic. Algebras under consideration have not to have a unit and have not to be associative. A multiplication on a vector space $W$ is a bilinear mapping $W\times W \to W$. We denote by $(W,P)$ the algebra with underlining space $W$ and multiplication $P$. Given a vector space $W$, a linear mapping $A:W\rightarrow W$, and a bilinear mapping $B:W\times W\to W$, we can define a multiplication $[A,B]:W\times W\to W$ by the formula
$$[A,B](x,y)=A(B(x,y))-B(A(x),y)-B(x,A(y))$$
for $x,y\in W$. For an algebra $A$ with a multiplication $P$ and $x\in A$ we denote by $L_x^P$ the operator of left multiplication by $x$. If the multiplication $P$ is fixed, we write $L_x$ instead of $L_x^P$.

In 1990 Kantor~\cite{Kantor90} defined the multiplication $\cdot$ on the set of all algebras (i.e. all multiplications) on the $n$-dimensional vector space $V_n$ as follows:
$$A\cdot B = [L_{e}^A,B],$$
where $A$ and $B$ are multiplications and $e\in V_n$ is some fixed vector. If $n>1$, then the algebra $W(n)$ does not belong to any well-known class of algebras (such as associative, Lie, Jordan, or Leibniz algebras). The algebra $W(n)$ turns out to be a conservative algebra (see below).

In 1972 Kantor~\cite{Kantor72} introduced conservative algebras as a generalization of Jordan algebras. Namely, an algebra $A=(W,P)$ is called a {\it conservative algebra} if there is a new multiplication $F:W\times W\rightarrow W$ such that
\begin{eqnarray}\label{tojd_oper}
[L_b^P, [L_a^P,P]]=-[L_{F(a,b)}^P,P]
\end{eqnarray}
for all $a,b\in W$. In other words, the following identity holds for all $a,b,x,y\in W$:
\begin{multline}\label{tojdestvo_glavnoe}
b(a(xy)-(ax)y-x(ay))-a((bx)y)+(a(bx))y+(bx)(ay)\\
-a(x(by))+(ax)(by)+x(a(by))
=-F(a,b)(xy)+(F(a,b)x)y+x(F(a,b)y).
\end{multline}
The algebra $(W,F)$ is called an algebra {\it associated} to $A$.

Let us recall some well-known results about conservative algebras. In~\cite{Kantor72} Kantor  classified all conservative 2-dimensional algebras and defined the class of {\it terminal} algebras as algebras satisfying some certain identity. He proved that every terminal algebra is a conservative algebra and classified all simple finite-dimensional terminal algebras with left quasi-unit over an algebraically closed field of characteristic zero~\cite{Kantor89term}. Terminal trilinear operations were studied in~\cite{Kantor89tril}, and some questions concerning classification of simple conservative algebras were considered in~\cite{Kantor89trudy}. After that, Cantarini and Kac classified simple finite-dimensional (and linearly compact) super-commutative and super-anticommutative conservative superalgebras and some generalization of these algebras (also known as ``rigid''{} superalgebras) over an algebraically closed field of characteristic zero \cite{kac_can_10}.

The algebra $W(n)$ plays a similar role in the theory of conservative algebras as the Lie algebra of all $n\times n$ matrices $gl_n$ plays in the theory of Lie algebras. Namely, in~\cite{Kantor88,Kantor90}  Kantor considered the category $\mathcal{S}_n$ whose objects are conservative algebras of non-Jacobi dimension $n$. It was proven that the algebra $W(n)$ is the universal attracting object in this category, i.e., for every $M\in\mathcal{S}_n$ there exists a canonical homomorphism from $M$ into the algebra $W(n)$. In particular, all Jordan algebras of dimension $n$ with unity are contained in the algebra
$W(n)$. The same statement also holds for all noncommutative Jordan algebras of dimension $n$ with unity.
Some properties of the product in the algebra $W(n)$ were studied in \cite{kay}.
The study of low dimensional conservative algebras was started in \cite{kaylopo},
where derivations and subalgebras of codimension 1 of the algebra $W(2)$ and of its principal subalgebras were described.

\section{Conservative algebra $W(2)$ and its subalgebras}

A multiplication on 2-dimensional vector space is defined by a $2\times 2\times 2$ matrix. Their  classification was given in many papers (see, for example,~\cite{GR11}).
Let consider the space $W(2)$ of all multiplications on the 2-dimensional space $V_2$ with a basis $v_1,v_2$.  The definition of the multiplication $\cdot$  on the algebra $W(2)$ can be found in  Section~1. Namely, we fix the vector $v_1 \in V_2$ and define
$$(A \cdot B)(x,y)=A(v_1,B(x,y))-B(A(v_1,x),y)-B(x,A(v_1,y))$$
for $x,y \in V_2$ and $A,B \in W(2)$. The algebra $W(2)$ is conservative \cite{Kantor90}.

Let consider the multiplications $\alpha_{i,j}^k$ ($i,j,k=1,2$) on $V_2$ defined by the formula $\alpha_{i,j}^k(v_t,v_l)=\delta_{it}\delta_{jl} v_k$ for all $t,l$. It is easy to see that $\{ \alpha_{i,j}^k | i,j,k=1,2 \}$ is a basis of the algebra $W(2)$. The multiplication tabel of $W(2)$ in this basis is given in \cite{kaylopo}. In this work we use another basis for the algebra $W(2)$. Let introduce the notation
$$
\begin{aligned}
e_1&=\alpha_{11}^1-\alpha_{12}^2-\alpha_{21}^2,
e_2=\alpha_{11}^2, e_3=\alpha_{22}^2-\alpha_{12}^1-\alpha_{21}^1,
e_4=\alpha_{22}^1, e_5=2\alpha_{11}^1+\alpha_{12}^2+\alpha_{21}^2,\\
e_6&=2\alpha_{22}^2+\alpha_{12}^1+\alpha_{21}^1,
e_7=\alpha_{12}^1-\alpha_{21}^1,
e_8=\alpha_{12}^2-\alpha_{21}^2.
\end{aligned}
$$
It is easy to see that the multiplication tabel of $W(2)$ in the basis $e_1,\dots,e_8$ is the following.

\begin{center}
\begin{tabular}{c|c|c|c|c|c|c|c|c|}
      & $e_1$ & $e_2$ & $e_3$ & $e_4$ & $e_5$ & $e_6$ & $e_7$ & $e_8$ \\ \hline
$e_1$ & $-e_1$ & $-3e_2$ & $e_3$ & $3e_4$ & $-e_5$ & $e_6$ & $e_7$ & $-e_8$ \\ \hline
$e_2$ & $3e_2$ & $0$ & $2e_1$ & $e_3$ & $0$ & $-e_5$ & $e_8$ & $0$ \\ \hline
$e_3$ & $-2e_3$ & $-e_1$ & $-3e_4$ & $0$ & $e_6$ & $0$ & $0$ & $-e_7$ \\ \hline
$e_4$ & $0$ & $0$ & $0$ & $0$ & $0$ & $0$ & $0$ & $0$ \\ \hline
$e_5$ & $-2e_1$ & $-3e_2$ & $-e_3$ & $0$ & $-2e_5$ & $-e_6$ & $-e_7$ & $-2e_8$ \\ \hline
$e_6$ & $2e_3$ & $e_1$ & $3e_4$ & $0$ & $-e_6$ & $0$ & $0$ & $e_7$ \\ \hline
$e_7$ & $2e_3$ & $e_1$ & $3e_4$ & $0$ & $-e_6$ & $0$ & $0$ & $e_7$ \\ \hline
$e_8$ & $0$ & $e_2$ & $-e_3$ & $-2e_4$ & $0$ & $-e_6$ & $-e_7$ & $0$ \\ \hline
\end{tabular}
\end{center}
\medskip
The subalgebra generated by the elements $e_1, \ldots, e_6$ is the conservative (and, moreover,  terminal) algebra $W_2$ of commutative 2-dimensional algebras. 
The subalgebra generated by the elements $e_1, \ldots, e_4$ is the conservative (and, moreover,  terminal) algebra $S_2$ of all commutative 2-dimensional algebras with trace zero multiplication \cite{kaylopo}.
We denote by $p_k:W(2)\rightarrow\FF$ ($k=1,\dots,8$) the map which sends $\sum\limits_{k=1}^8a_ke_k$ to $a_k$.

\section{One-sided ideals}

\subsection{One-sided ideals in $W(2)$} 
Let $Ann_l(W(2))$ be the space of $W(2)$ generated by the elements
$e_4, e_5-2e_1-3e_8, e_3+e_6$, and $e_3+e_7.$
It is easy to see that $xW(2)=0$ for any $x \in Ann_l(W(2))$.

\begin{Lem}\label{righideal}
If $I$ is a non-trivial right ideal of $W(2)$,
then $I$ is a subspace of $Ann_l(W(2)).$
\end{Lem}

{\bf Proof.}
Let $x\in I.$ If $p_2(x) \neq 0$, then
$(xe_4)e_2 = -p_2(x) e_1 \in I.$
It follows from the multiplication table of the algebra $W(2)$ that $I=W(2)$.
Let now consider the case where $p_2(x)=0$  for any $x\in I$.

We have $(x e_3)e_2=- (p_1-p_5-p_8)(x)e_1 \in I$.
So $I=W(2)$ or $p_1(x)=p_5(x)+p_8(x)$ for any $x\in I$.
We have
$$xe_2= (-p_3+p_6+p_7)(x)e_1+(-3p_1-3p_5+p_8)(x)e_2.$$
Then
$p_8(x)=3(p_1+p_5)(x) \mbox{ and } p_3(x)=(p_6+p_7)(x)$.
So any element $x\in I$ has the form
$$x=p_5(x)(e_5-2e_1-3e_8)+p_4(x)e_4+p_6(x)(e_3+e_6)+p_7(x)(e_3+e_7)\in Ann_l(W(2)).$$
The lemma is proved.

\begin{Lem}\label{leftideal}
If $I$ is a non-trivial left  ideal of $W(2)$,
then $I$ is one of the following spaces:

1) $W_{\alpha,\beta}={}_{\FF}\langle e_1,e_2,e_3,e_4,  \alpha e_5 +\beta e_8, -\alpha e_6 + \beta e_7\rangle$, where
$\alpha, \beta \in \FF$, $(\alpha,\beta)\not=(0,0)$;

2) $I_1={}_{\FF}\langle e_1,e_2,e_3,e_4\rangle;$

3) $I_2={}_{\FF}\langle e_5,e_6,e_7,e_8\rangle;$

4) $w_{\alpha,\beta}={}_{\FF}\langle\alpha e_5 +\beta e_8, -\alpha e_6 + \beta e_7\rangle,$
where $\alpha, \beta \in \FF$, $(\alpha,\beta)\not=(0,0)$.

\end{Lem}

{\bf Proof.}
Let $x=\in I.$ We have
$e_2(e_2(e_1x-x))= 4 p_4(x) e_1 \in I.$

{\bf 1.} Assume that there is $x\in I$ such that $p_4(x) \neq 0.$ 
Then it is obvious that $e_1,e_2,e_3,e_4 \in I.$
In this case either $I$ is the subspace generated by $e_1,e_2,e_3,e_4$ (in this case $I=I_1$) or 
there exists a nonzero element 
$w=\sum\limits_{i=5}^8 \alpha_i e_i  \in I.$
In the former case we have
\begin{eqnarray*}
 w_1=w-e_1w &=& 2\alpha_5e_5+2\alpha_8e_8\in I,\\ 
 w_2=e_2(e_1w+w)&=&-2\alpha_6e_5+2\alpha_7e_8\in I,\\
 w_3=e_1w+w &=& 2\alpha_6e_6+2\alpha_7e_7 \in I,\\
 w_4=e_2(w-e_1w)&=&-2\alpha_5e_6+2\alpha_8e_7 \in I.\\
 \end{eqnarray*}
Now, if elements $w_1$ and $w_2$ are linear independent, then  $e_5,e_6,e_7,e_8 \in I$ and $I=W(2).$
If the elements  $w_1$ and $w_2$ are linear dependent, then $I=W_{\alpha,\beta}$ for some $\alpha,\beta\in\FF$.

{\bf 2.} Assume that $p_4(x)=0$ for any $x\in I$.
If $p_i(x)=0$ for any $x\in I$ and $i=1,2,3,4,$
then by the previous argumention, we have either $I=I_2$ or $I=w_{\alpha,\beta}$ for some $\alpha,\beta\in\FF$.

Let consider $x\in I$. If $p_1(x)\neq 0$, then we have $$e_6(e_3(e_3(e_2(e_1x-x))))=-36 p_1(x) e_4 \in I$$
and we have a contradiction with $p_4(I)=0.$

Suppose that $p_1(I)=p_4(I)=0$. 
If $p_2(x)\neq 0$ or $p_3(x)\neq0$ for some $x\in I$, then
we have $$e_3(e_1x+x)=2p_2(x)e_1 -6p_3(x) e_4 \in I$$
and we have a contradiction with $p_1(I)=p_4(I)=0$.
The lemma is proved.

\subsection{One-sided ideals in $W_2$}
Let $Ann_l(W_2)$ be the subspace of $W_2$ generated by the elements
$e_4$, and $e_3+e_6.$
It is easy to see that $xW_2=0$ for any $x \in Ann_l(W_2)$.

\begin{Lem}
If $I$ is a non-trivial right ideal of $W_2$,
then $I$ is a subspace of $Ann_l(W_2).$
\end{Lem}

{\bf Proof.} The proof is analogous to the proof of Lemma \ref{righideal}.

\begin{Lem}\label{leftW2}
If $I$ is a non-trivial left  ideal of $W_2$,
then $I$ is one of the following spaces:

1) ${}_{\FF}\langle e_1,e_2,e_3,e_4\rangle;$

2) ${}_{\FF}\langle e_5,e_6\rangle.$

\end{Lem}

{\bf Proof.} The proof is analogous to the proof of Lemma \ref{leftideal}.

\subsection{One-sided ideals in $S_2$}
Let $Ann_l(S_2)$ be the subspace of $S_2$ generated by the  element
$e_4.$
It is easy to see that $xS_2=0$ for any $x \in Ann_l(S_2)$.

\begin{Lem}
If $I$ is a non-trivial right ideal of $S_2$, then $I$ is a subspace of $Ann_l(S_2).$
\end{Lem}

{\bf Proof.} The proof is analogous to the proof of Lemma \ref{righideal}.

\begin{Lem}
There are no non-trivial left ideals in $S_2$.
\end{Lem}

{\bf Proof.} The proof is analogous to the proof of Lemma \ref{leftideal}.

\subsection{Corollaries} 
Here we want to formulate some corollaries about a decomposition of the algebras under consideration in a sum of one-sided ideals and about their ternary derivations.

\begin{corollary} For any left ideal $Y_1$ of the algebra $W \in \{W(2), W_2,S_2\}$ there is some left ideal $Y_2$
of $W$ such that $W=Y_1+Y_2$ and $Y_1 \cap Y_2=0.$
\end{corollary}

The definition of a ternary derivation arised in the paper of Perez-Izquierdo and Jimenez-Gestal \cite{perez03}.
A ternary derivation is a triple of linear mappings $(D,F,G)$ such that $D(xy)=F(x)y+xG(y).$
A ternary derivation is trivial if $D$, $F$, and $G$ are sums of a derivation and an element of a centroid of an algebra under consideration. Ternary derivations of Jordan algebras and superalgebras, and $n$-ary algebras were studied in \cite{kay13,kay14,shestakov14}.

\begin{corollary} The algebras $W(2), W_2$ and $S_2$ have non-trivial ternary derivations.
\end{corollary}

{\bf Proof.} Let $A$ be an algebra from the set $\{W(2),W_2,S_2\}.$
By the lemmas above $A$ has a nonzero left annihilator $Ann_l(A)$.
Then for any nonzero linear map $\phi$ such that $\phi(A) \subseteq Ann_l(A)$  the triple $(0, \phi, 0)$ is a non-trivial ternary derivation of $A$.
The corollary is proved.

\medskip

\section{Automorphisms}

In this section we discuss the automorphism groups of the algebras $W(2)$, $W_2$ and $S_2$. Firstly we prove a lemma which provides a series of automorphisms of the algebra $W(n)$ for any $n\ge 2$. Let fix some space $V$ and nonzero element $e\in V$. We denote by $W(V,e)$ the algebra whose elements are bilinear maps from $V\times V$ to $V$ and multiplication on which is defined by the formula
$$
(A\cdot B)(x,y)=A(e,B(x,y))-B(A(e,x),y)-B(x,A(e,y))
$$
for $A,B\in W(V,e)$, $x,y\in V$. Let introduce the notation $GL(V,e)=\{f\in GL(V)\mid f(e)=e\}$. It is clear that $GL(V,e)$ is a subgroup of $GL(V)$. If $V$ is an $n$-dimensional vector space, then $GL(V,e)$ is isomorphic to $(n-1)$-dimensional {\it affine group} over $\FF$.

\begin{Lem}\label{auttoaut}
There is a monomorphism of groups $\Phi:GL(V,e)\rightarrow \Aut\big(W(V,e)\big)$ defined by the equality
$$
\big(\Phi(f)(A)\big)(x,y)=fA(f^{-1}(x),f^{-1}(y))
$$
for $f\in GL(V,e)$, $A\in W(V,e)$, $x,y\in V.$
\end{Lem}
{\bf Proof.}
Let us prove that $\Phi(f)\in \Aut\big(W(V,e)\big)$ for any $f\in GL(V,e)$. Actually, for $A,B\in W(V,e)$, $x,y\in V$ we have
\begin{multline*}
\big(\Phi(f)(A)\cdot \Phi(f)(B)\big)(x,y)=
fA\big(f^{-1}(e),f^{-1}fB(f^{-1}(x),f^{-1}(y))\big)\\
-fB\big(f^{-1}fA(f^{-1}(e),f^{-1}(x)),f^{-1}(y)\big)
-fB\big(f^{-1}(x),f^{-1}fA(f^{-1}(e),f^{-1}(y))\big)\\
=f\Big(A\big(e,B(f^{-1}(x),f^{-1}(y))\big)
-B\big(A(e,f^{-1}(x)),f^{-1}(y)\big)
-B\big(f^{-1}(x),A(e,f^{-1}(y))\big)\Big)\\
=\big(\Phi(f)(A\cdot B)\big)(x,y).
\end{multline*}
It is easy to check that $\Phi$ is a homomorphism of groups.
Let $f\in GL(V,e)$ be such that $f\not=\Id_V$. Then there is such $x\in V$ that $f^{-1}(x)\not=x$. It is clear that there is such $A\in W(V,e)$ that $A(e,x)=e$ and $A(e,f^{-1}(x))\not=e$. Then
$$\Phi(A)(e,x)=fA(f^{-1}(e),f^{-1}(x))=fA(e,f^{-1}(x))\not=f(e)=A(e,x).$$
So $\Phi(f)\not=\Id_{W(V,e)}$ and we prove that $\Phi$ is a monomorphism.
The Lemma is proved.

\medskip

Any element of the group $GL(V_2,v_1)$ can be written as a composition $\SSS_b\TT_a=\TT_{ab}\SSS_b$ ($a,b\in\FF$, $b\not=0$), where $\TT_a(v_2)=v_2+av_1$ and $\SSS_b(v_2)=\frac{v_2}{b}$.
We write simply $\TT_a$ and $\SSS_b$ instead of $\Phi(\TT_a)$ and $\Phi(\SSS_b)$. Direct calculations show that the action of $\TT_a$ and $\SSS_b$ on $W(V_2)$ is defined by the formulas
\begin{multline*}
\TT_a(e_1)=e_1+2ae_3+3a^2e_4,\:\:\TT_a(e_2)=e_2+ae_1+a^2e_3+a^3e_4,\:\:\TT_a(e_3)=e_3+3ae_4,\:\:\TT_a(e_4)=e_4,\\
\TT_a(e_5)=e_5-ae_6,\:\:\TT_a(e_6)=e_6,\:\:\TT_a(e_7)=e_7,\:\:\TT_a(e_8)=e_8+ae_7
\end{multline*}
and
\begin{multline*}
\SSS_b(e_1)=e_1,\:\:\SSS_b(e_2)=\frac{e_2}{b},\:\:
\SSS_b(e_3)=be_3,\:\:\SSS_b(e_4)=b^2e_4,\\
\SSS_b(e_5)=e_5,\:\:\SSS_b(e_6)=be_6,\:\:
\SSS_b(e_7)=be_7,\:\:\SSS_b(e_8)=e_8.
\end{multline*}
It follows from these formulas that $\TT_a$ and $\SSS_b$ induce automorphisms of $W_2$ and $\SSS_2$ which we denote by $\TT_a$ and $\SSS_b$ too.

\begin{Remark}
It follows from the proof of \cite[Theorem 3]{kaylopo} that $L_{e_7}$ and $L_{e_8}$ are derivations of the algebra $W(2)$. The map $L_{e_7}$ is nilpotent and it is easy to see that $\TT_a=\sum_{k=0}^{\infty}\frac{(aL_{e_7})^k}{k!}$.
On the other hand the map $L_{e_8}$ is not nilpotent. But if the element $e^t=\sum_{k=0}^{\infty}\frac{t^k}{k!}$ is defined and lies in $\FF^*$ for some $t\in\FF$, then we have an equality $\SSS_{e^t}=\sum_{k=0}^{\infty}\frac{(tL_{e_8})^k}{k!}$. So if for any $b\in\FF^*$ there is $t\in\FF$ such that $b=e^t$, then the automorphism $\SSS_b$ can be defined using  the derivation $L_{e_8}$.
\end{Remark}

Now we are ready to describe the automorphism groups of the algebras $W(2)$, $W_2$ and $S_2$.

\begin{Th}
The homomorphism $\Phi:GL(V_2,v_1)\rightarrow \Aut\big(W(2)\big)$ is bijective. Moreover, the induced homomorphisms from $GL(V_2,v_1)$ to $\Aut(W_2)$ and $\Aut(S_2)$ are bijective too. In other words, $\Aut\big(W(2)\big)\cong \Aut(W_2)\cong \Aut(S_2)$ is isomorphic to the matrix group $\begin{pmatrix}1 & \FF\\ 0 & \FF^*\end{pmatrix}$ (the one-dimensional affine group over $\FF$).
\end{Th}
{\bf Proof.} It can be easily checked that all three homomorphisms of groups considered in the theorem are monomorphic. So it is enough to prove that they are epimorphic. Firstly we prove that the automorphisms $\TT_a$ ($a\in\FF$) and $\SSS_b$ ($b\in\FF^*$) generate the group $\Aut(S_2)$.

Let us take some $f\in\Aut(S_2)$. Since $f$ sends any right ideal to a right ideal, we have $f(e_4)=te_4$ for some $t\in\FF^*$. Then we have $0=f(e_3e_4)=tf(e_3)e_4$. So $f(e_3)=be_3+ae_4$ for some $a,b\in\FF$. Let us consider the equality $-3f(e_4)=f(e_3)f(e_3)$. It can be rewritten in the form $-3te_4=-3b^2e_4$. So $t=b^2$. Let us consider $g=\SSS_{\frac{1}{b}}\TT_{-\frac{a}{3b}}f$. We know that $g(e_3)=e_3$ and $g(e_4)=e_4$ and it remains to prove that $g=\Id_{S_2}$. It follows from the equality $e_3=g(e_1e_3)=g(e_1)e_3$ that $g(e_1)=e_1+te_4$ for some $t\in\FF$. Then we have $$-e_1-te_4=g(-e_1)=g(e_1e_1)=\big(g(e_1)\big)^2=-e_1+3te_4.$$
So $t=0$ and we have $g(e_1)=e_1$. Using the equality $2e_1=g(e_2e_3)=g(e_2)e_3$ we obtain that $g(e_2)=e_2+te_4$ for some $t\in\FF$. Consideration of the equality $0=\big(g(e_2)\big)^2$ shows that $t=0$ and $g(e_2)=e_2$. So we prove that $\Aut(S_2)$ is generated by $\TT_a$ and $\SSS_b$ ($a,b\in\FF$, $b\not=0$).

Let now prove that $\TT_a$ ($a\in\FF$) and $\SSS_b$ ($b\in\FF^*$) generate the group $\Aut(W_2)$. Since $S_2$ is the unique 4-dimensional left ideal of $W_2$ by Lemma \ref{leftW2}, every automorphism of $W_2$ induces an automorphism of $S_2$. Let take some $f\in\Aut(W_2)$. Since $\Aut(S_2)$ is generated by automorphisms of the form $\TT_a$ and $\SSS_b$, there are such $a\in\FF$ and $b\in\FF^*$ that the automorphism $g=\SSS_b\TT_af$ is identical on $S_2$. It remains to prove that $g(e_5)=e_5$ and $g(e_6)=e_6$.
But $e_5$ and $e_6$ generate the unique 2-dimensional left ideal of $W_2$ and so $g(e_5)=t_1e_5+t_2e_6$, $g(e_6)=t_3e_5+t_4e_6$. Consideration of the equalities $-3e_2=g(e_5e_2)=g(e_5)e_2$ and $e_1=g(e_6e_2)=g(e_6)e_2$ shows that $t_1=t_4=1$ and $t_2=t_3=0$, i.e. $g=\Id_{W_2}$.

It remains to prove that $\TT_a$ ($a\in\FF$) and $\SSS_b$ ($b\in\FF^*$) generate the group $\Aut\big(W(2)\big)$. By Lemma \ref{leftideal} the algebra $W(2)$ has two 4-dimensional left ideals: $S_2$ and $I={}_{\FF}\langle e_5,e_6,e_7,e_8\rangle$. Note that $W(2)$ contains 2-dimensional ideals and they all lie in $I$. So any automorphism of the algebra $W(2)$ maps $S_2$ to $S_2$ and $I$ to $I$. Analogously to the case of the algebra $W_2$ it is enough to prove that any automorphism $g\in\Aut\big(W(2)\big)$ identical on $S_2$ is identical on $I$. It follows from the equalities $g(e_5)=-e_1g(e_5)$ and $g(e_8)=-e_1g(e_8)$ that the space ${}_{\FF}\langle e_5,e_8\rangle$ is invariant under the action of $g$. Using the equalities $-2e_1=g(e_5)e_1$ and $0=g(e_5)e_4$ we see that $g(e_5)=e_5$. Using the equalities $0=g(e_8)e_1$ and $-2e_4=g(e_8)e_4$ we see that $g(e_8)=e_8$. Finally, we have $g(e_6)=g(e_3e_5)=e_3e_5=e_6$ and $g(e_7)=-g(e_3e_8)=-e_3e_8=e_7$. So the Theorem is proved.

\begin{corollary}\label{isoaut}
Let $A$ and $B$ be two multiplications on the space $V_2$. There is an isomorphism from the algebra $(V_2,A)$ to the algebra $(V_2,B)$ which sends $v_1$ to $v_1$ iff there is an automorphism of the algebra $W(2)$ which sends $(V_2,A)$ to $(V_2,B)$.
\end{corollary}

\section{Idempotents}

Let now describe the idempotents of the algebra $W(2)$. For $e\in W(2)$ we denote by $\OO(e)=\{\SSS_b\TT_a(e)\mid a,b\in\FF,b\not=0\}$ the orbit of $e$ under the action of the automorphism group of $W(2)$. It is clear that $e$ is an idempotent iff all elements of $\OO(e)$ are idempotents.
We denote by $\bar\FF$ the set of representatives of cosets of the subgroup $(\FF^*)^2$ in the group $\FF^*$.

\begin{Th}\label{idemp}
The set of nonzero idempotents of the algebra $W(2)$ equals to the disjoint union of the following sets:
\begin{enumerate}
\item $\OO(e_8+e_2-e_1+c(3e_8+e_5-2e_1))$ ($c\in\FF$);
\item $\OO(-e_1+c(e_5-2e_1)+de_8)$ ($c,d\in\FF$);
\item $\OO(-e_1-2e_8+4e_3+e_6+3e_7+c(3e_8-e_5+2e_1)+de_4)$ ($c,d\in\FF$);
\item $\OO(-e_1-2e_8+c(3e_8-e_5+2e_1)+qe_4)$ ($c\in\FF$, $q\in\bar\FF$);
\end{enumerate}
\end{Th}
{\bf Proof.} Direct calculations show that all elements of the set described in the theorem are idempotent. Let $e\in W(2)$ be a nonzero idempotent. Let us prove that $e$ lies in the set described in the theorem. For $c,d\in \FF$, $q\in\bar\FF$ let introduce the notation
\begin{multline*}
w_1(c)=e_8+e_2-e_1+c(3e_8+e_5-2e_1),
w_2(c,d)=-e_1+c(e_5-2e_1)+de_8,\\
w_3(c,d)=-e_1-2e_8+4e_3+e_6+3e_7+c(3e_8-e_5+2e_1)+de_4,\\
w_4(c,d)=-e_1-2e_8+c(3e_8-e_5+2e_1)+de_4.
\end{multline*}

{\bf 1.} Assume that $p_2(e)\not=0$. Let us take
$b=p_2(e)$ and $a=-\frac{p_1(e)+2p_5(e)}{p_2(e)}-1$.
It is easy to see that $p_2\SSS_b\TT_a(e)=1$ and $(p_1+2p_5)\SSS_b\TT_a(e)=-1$. So we may suppose that $p_2(e)=1$ and $(p_1+2p_5)(e)=-1$ initially.
Let us introduce the notation $c=p_5(e)$. Then we have $p_1(e)=-(2c+1)$. It follows from the equality $p_2(e^2)=p_2(e)$ that $p_8(e)=3c+1$. Rewriting the equalities $p_5(e^2)=p_5(e)$ and $p_8(e^2)=p_8(e)$ we obtain that $p_6(e)=p_7(e)=0$. Rewriting the equality $p_1(e^2)=p_1(e)$ we obtain $p_3(e)=-p_6(e)-p_7(e)=0$. Then we have $0=p_3(e^2)=p_2(e)p_4(e)=p_4(e)$. So we have
$$
e=\sum\limits_{k=1}^8p_k(e)e_k=-(2c+1)e_1+e_2+ce_5+(3c+1)e_8=w_1(c).
$$

Let now prove that $(p_1+2p_5)(e)=-1$ if $e$ is a nonzero idempotent such that $p_2(e)=0$. Using multiplication table we see that
$$p_1(e^2)=-\big(p_1(e)+2p_5(e)\big)p_1(e),\:\:
p_5(e^2)=-\big(p_1(e)+2p_5(e)\big)p_5(e).$$
Since $e^2=e$, we have $p_1(e)+2p_5(e)=-1$ in the case where $p_1(e)\not=0$ or $p_5(e)\not=0$. It is enough to prove that $e=0$ if $p_1(e)=p_2(e)=p_5(e)=0$. But in this case we obtain consequently using multiplication table and the equality $e^2=e$  that $p_8(e)=0$, $p_6(e)=p_7(e)=0$, $p_3(e)=0$ and $p_4(e)=0$, i.e. $e=0$.

Also, if $p_2(e)=0$, we have the equalities
$$
\begin{aligned}
p_3(e)&=p_3(e^2)=2(p_6(e)+p_7(e)-p_3(e))p_1(e)-(p_8(e)+p_5(e)-p_1(e))p_3(e),\\
p_4(e)&=p_4(e^2)=3(p_6(e)+p_7(e)-p_3(e))p_3(e)-(2p_8(e)-3p_1(e))p_4(e),\\
p_6(e)&=p_6(e^2)=(p_3(e)-p_6(e)-p_7(e))p_5(e)-(p_8(e)+p_5(e)-p_1(e))p_6(e),\\
p_7(e)&=p_7(e^2)=(p_6(e)+p_7(e)-p_3(e))p_8(e)-(p_8(e)+p_5(e)-p_1(e))p_7(e),
\end{aligned}
$$
which can be rewritten in the form
\begin{equation}\label{eqs}
\begin{aligned}
(p_8(e)+p_5(e)-p_1(e)+1)p_3(e)&=2(p_6(e)+p_7(e)-p_3(e))p_1(e),\\
(2p_8(e)-3p_1(e)+1)p_4(e)&=3(p_6(e)+p_7(e)-p_3(e))p_3(e),\\
(p_8(e)+p_5(e)-p_1(e)+1)p_6(e)&=(p_3(e)-p_6(e)-p_7(e))p_5(e),\\
(p_8(e)+p_5(e)-p_1(e)+1)p_7(e)&=(p_6(e)+p_7(e)-p_3(e))p_8(e).
\end{aligned}
\end{equation}

{\bf 2.} Assume that $p_2(e)=0$, $p_8(e)-p_5(e)-2p_1(e)\not=0$. Let us take
$a=-\frac{p_3(e)-p_6(e)-p_7(e)}{p_8(e)-p_5(e)-2p_1(e)}$.
It is easy to see that $(p_3-p_6-p_7)\TT_a(e)=0$. So we may suppose that $(p_3-p_6-p_7)(e)=0$ initially.
Since $$2p_8(e)-3p_1(e)+1=2(p_8(e)+p_5(e)-p_1(e)+1)=2(p_8(e)+3p_5(e))\not=0,$$ it follows from \eqref{eqs} that $p_3(e)=p_4(e)=p_6(e)=p_7(e)=0$. Let us denote $p_5(e)$ by $c$ and $p_8(e)$ by $d$. Then $p_1(e)=-(2c+1)$ and we have
$e=w_2(c,d)$.

{\bf 3.} Assume that $p_2(e)=0$, $p_8(e)-p_5(e)-2p_1(e)=0$. If $p_6(e)+p_7(e)-p_3(e)\not=0$, then it follows from \eqref{eqs} that $p_1(e)=p_5(e)=0$. This contradicts the equality $p_1(e)+2p_5(e)=-1$ proved above. So we have $p_3(e)=p_6(e)+p_7(e)$. Let us take $a=\frac{p_3(e)-4p_6(e)}{2}$. It is easy to see that $(4p_6-p_3)\TT_a(e)=0$. So we may suppose that $(4p_6-p_3)(e)=0$ initially. Let now consider two cases.

{\bf 3a.} Let $p_6(e)\not=0$. Then we have $p_6\SSS_{\frac{1}{p_6(e)}}(e)=1$ and so we may suppose that $p_6(e)=1$ initially. Then we have $p_3(e)=4$ and $p_7(e)=p_3(e)-p_6(e)=3$. Let us denote $p_5(e)$ by $-c$ and $p_4(e)$ by $d$. We have $p_1(e)=2c-1$ and $p_8(e)=p_5(e)+2p_1(e)=3c-2$. So
$e=w_3(c,d)$.

{\bf 3b.} Let $p_6(e)=0$. Then we have $p_3(e)=p_7(e)=0$. If $p_4(e)=0$, then we have
$e=w_2(c,d)$ for $c=p_5(e)$ and $d=p_8(e)$. Let now $p_4(e)\not=0$. There are $b\in\FF^*$ and $q\in\bar\FF$ such that $b^2p_4(e)=q$. If we denote $p_5(e)$ by $-c$, we obtain
$\SSS_b(e)=w_4(c,q)$.

Let us denote by $E$ the set $$\{w_1(c)\}_{c\in\FF}\cup\{w_2(c,d)\}_{c,d\in\FF}\cup \{w_3(c,d)\}_{c,d\in\FF}\cup\{w_4(c,q)\}_{c\in\FF, q\in\bar\FF}.$$
We prove that for any $w\in E$ and $f\in\Aut\big(W(2)\big)$ the condition $f(w)\in E$ is equivalent to $f(w)=w$ (it is clear that the union of sets listed in the theorem is disjoint in this case).

{\bf 1.} Suppose that $e=\SSS_b\TT_a(w_1(c))\in E$ ($a,c\in\FF$, $b\in\FF^*$). Since $p_2(e)=\frac{1}{b}$ and $p_2(w)\in\{0,1\}$ for any $w\in E$, we have $b=1$. Then we have $(p_1+2p_5)(e)=a-1$. Since $e\in E$ and $p_2(e)=1$, we have $(p_1+2p_5)(e)=-1$ and so $a=0$. We obtain that $e=w_1(c)$.

{\bf 2.} Suppose that $e=\SSS_b\TT_a(w_2(c,d))\in E$ ($a,c,d\in\FF$, $b\in\FF^*$). We have $p_3(e)=-4abc-2ab$, $p_6(e)=-abc$. We have $p_3(w)=4p_6(w)$ for any $w\in E$. Since $e\in E$, we have $-4abc-2ab=-4abc$, i.e. $a=0$. Then $e=\SSS_b(w_2(c,d))=w_2(c,d)$.

{\bf 3.} Suppose that $e=\SSS_b\TT_a(w_3(c,d))\in E$ ($a,c,d\in\FF$, $b\in\FF^*$). Analogously to the previous point we have $a=0$. Then $p_6(e)=b$. Since $p_6(w)\in\{0,1\}$ for any $w\in E$, we have $b=1$. Then $e=w_3(c,d)$.

{\bf 4.} Suppose that $e=\SSS_b\TT_a(w_4(c,q))\in E$ ($a,c\in\FF$, $b\in\FF^*$, $q\in\bar\FF$). Analogously to the second point of the present proof we have $a=0$. Then $e=-e_1-2e_8+c(3e_8-e_5+2e_1)+b^2qe_4$.
So $e\in E$ iff $b^2q=0$ or $b^2q\in\bar\FF$. Since $b\not=0$ and $q\not=0$, $b^2q\in\bar\FF$. But $q$ and $b^2q$ lie in the same coset of the group $(F^*)^2$ in the group $F^*$ and so by definition only one of them lies in $\bar\FF$ if $q\not=b^2q$.
Since $q$ lies in $\bar\FF$, we have $b^2q=q$ and so $e=-e_1-2e_8+c(3e_8-e_5+2e_1)+qe_4=w_4(c,q)$.
The theorem is proved.

\medskip

Let introduce the algebras $W_1(t)$, $W_2(t,s)$, $W_3(t,s)$ and $W_4(t,u)$ for $s,t\in\FF$, $u\in\bar\FF$ by the following multiplication tables

\begin{center}
\begin{tabular}{llcll}
$W_1(t):$&
\begin{tabular}{c|c|c|}
&$v_1$&$v_2$\\ \hline
$v_1$&$-v_1+v_2$&$tv_2$\\ \hline
$v_2$&$0$&$0$\\ \hline
\end{tabular}&\phantom{xxxx}&
$W_2(t,s)$:&
\begin{tabular}{c|c|c|}
&$v_1$&$v_2$\\ \hline
$v_1$&$-v_1$&$tv_2$\\ \hline
$v_2$&$sv_2$&$0$\\ \hline
\end{tabular}\\
\\
$W_3(t,s)$:&
\begin{tabular}{c|c|c|}
&$v_1$&$v_2$\\ \hline
$v_1$&$-v_1$&$-v_2$\\ \hline
$v_2$&$tv_2-v_1$&$v_2+sv_1$\\ \hline
\end{tabular}&\phantom{xxxx}&
$W_4(t,u)$:&
\begin{tabular}{c|c|c|}
&$v_1$&$v_2$\\ \hline
$v_1$&$-v_1$&$-v_2$\\ \hline
$v_2$&$tv_2$&$uv_1$\\ \hline
\end{tabular}
\end{tabular}
\end{center}

It is easy to see that the algebra $W_1(t)$ corresponds to the element $w_1\left(\frac{t-2}{6}\right)\in W(2)$, the algebra $W_2(t,s)$ corresponds to the element $w_2\left(\frac{t+s-2}{6},\frac{t-s}{2}\right)\in W(2)$, the algebra $W_3(t,s)$ corresponds to the element $\SSS_{\frac{1}{6}}\left(w_3\left(\frac{3-t}{6},36s\right)\right)\in W(2)$, and the algebra $W_4(t,q)$ corresponds to the element $w_4\left(\frac{3-t}{6},q\right)\in W(2)$ (see the proof of Theorem \ref{idemp} for notation). We denote by $\mathcal{L}$ the set
$$\{W_1(t)\}_{t\in\FF}\cup\{W_2(t,s)\}_{t,s\in\FF}\cup \{W_3(t,s)\}_{t,s\in\FF}\cup\{W_4(t,u)\}_{t\in\FF, u\in\bar\FF}.$$

\begin{corollary}
Let $R$ be a nonzero two-dimensional algebra with left quasi-unit $e$. Then there is an unique algebra $W\in\mathcal{L}$ such that there is an isomorphism $\phi:R\rightarrow W$ satisfiing the condition $\phi(e)=v_1$.
\end{corollary}

{\bf Proof.} Let $\phi_0:R\rightarrow V_2$ be some isomorphism of vector spaces such that $\phi_0(e)=v_1$. There is an unique multiplication $A$ on $V_2$ such that $\phi_0$ is an isomorphism from $R$ to $(V_2,A)$. Then $v_1$ is a quasi-unit of $(V_2,A)$ and so $(V_2,A)$ is an idempotent of the algebra $W(2)$. By Theorem \ref{idemp} and the definition of the set $\mathcal{L}$ there are such $W\in\mathcal{L}$ and automorphism $f$ of $W(2)$ that $f$ sends the element of $W(2)$ corresponding $(V_2,A)$ to the element of $W(2)$ corresponding $W$. Then by Corollary \ref{isoaut} there is an isomorphism $\phi_1:(V_2,A)\rightarrow W$ such that $\phi_1(v_1)=v_1$. The map $\phi=\phi_1\phi_0$ satisfies to all required conditions. If we have some another algebra $W'\in\mathcal{L}$ and isomorphism $\phi':R\rightarrow W'$ such that $\phi'(e)=v_1$, then there is an isomorphism $\phi'\phi^{-1}:W\rightarrow W'$ which sends $v_1$ to $v_1$. By Corollary \ref{isoaut} the elements of $W(2)$ corresponding to $W$ and $W'$ lie in the same orbit under the action of $\Aut\big(W(2)\big)$. But it follows from definition of $\mathcal{L}$ and Theorem \ref{idemp} that $W=W'$ in this case.

\medskip 

{\bf Acknowledgements:}
The  first author was supported by the Brazilian FAPESP, Proc. 2014/24519-8,
the second author was supported by the Brazilian FAPESP, Proc. 2014/19521-3,
all authors were supported by RFBR 15-31-21169.
We are grateful for the support.

\end{document}